OpenStudy (anonymous):

Select all the expressions that are equivalent to 6y-5x 2(3y)-5x -5x+6y 5x-6y 6y-15/3x 1xy

TheSmartOne (thesmartone):

Do you have any idea?

OpenStudy (alexandervonhumboldt2):

hint 6=2*3 15/3=5

OpenStudy (alexandervonhumboldt2):

a-b=-b+a

TheSmartOne (thesmartone):

Besides those hints: a - b = -b + a

TheSmartOne (thesmartone):

Alex beat me to it x'D

OpenStudy (alexandervonhumboldt2):

yes i did haha

OpenStudy (anonymous):

is it A and D

OpenStudy (alexandervonhumboldt2):

A and D is correct, but there is 1 more

OpenStudy (alexandervonhumboldt2):

use ruel a-b=-b+a

OpenStudy (alexandervonhumboldt2):

thus 6y-5x=?

OpenStudy (anonymous):

i think its -5x + 6y

TheSmartOne (thesmartone):

and that's your final answer :)

OpenStudy (anonymous):

oh ok i get it now thank you

TheSmartOne (thesmartone):

No problem :)